Explain how increases in C, I, or G, or decreases in T cause output to increase by more than the initial increase in AE ?
1
Expert's answer
2017-06-18T04:13:10-0400
A multiplied effect of increase in C, I, or G, or decreases in T on total output takes place because consumption depends on total income (or total output). When the last one increases in response to mentioned changes in C, I, T and G, consumption also enlarges and income rises further. This chain effect is infinite and depends on MPC - marginal propensity to consume.
These multipliers are derived as infinite geometric series:
∆Y/∆G = ∆Y/∆I = 1/ (1 - MPC)
∆Y/∆T = -MPC/ (1 - MPC).
∆Y/∆C = 1/MPC